- T0290403522 was too low, TCS can't guide (not even manually.) Lost pointing, tel beyond limits. J. Rojas recovered by moving the RA drive by hand. One comm cable, at the basis of the telescope, now looks damaged, almost cut.
- NOS problems w/lines 15 and 16 (and others....). Tried ThAr lamp by hand to replace line 16, system freezes in "setting detector binning". Rebooted chiron twice. BUT exposure # 18 fails to start, "pan expose" did not help either and NOS is frozen again (re send did not reflect in the TCS) mode UNDEF and cannot set by hand (system freezes on "setting detector size".) Rebooted chiron once again (also, it was rebooted in the afternoon.) Then exposure 18 starts but "exposure time" stops in 11.64 seconds while in 'panview debug chiron' the time was still running and after a while the exposure time finally catches up. NOS line # 19 fails to start! And again it did not work to send it from OPTGUI, rebooted optgui once again and then it was possible to run the ThAr lamp. But NOS line #21 do not start, optgui mode undef. New reboot of optgui this time it reports "wait port 5325".
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-01-14 | 6.5 | 0.0 | 2.0 | 0.0 | 0.0 | 8.5 |
- Tololo's DIMM still down, used Pachon seeing monitor as a reference only.
- Lost chance to observe: HIP11766, HD223807, HIP000065.
- File 1139 complements files 34 and 35 (my glitch.)
- Added observations of LHS0288.
- Completed last integration (30 + 5 integrations) on Alpha Cen B at 08:59 UT, with the Sun at -11.8 degrees of altitude.